Sarah test average 93 and Sam test average is 87 in 7 tests.they both got 100 on 8 test.what is their new average
Answer by lwsshak3(11628)

Sarah test average 93 and Sam test average is 87 in 7 tests.they both got 100 on 8 test.what is their new average
***
Sarah's new average=(7*93+100)/8=751/8≈93.9
Sam's new average=(7*87+100)/8=709/8≈88.6